David borrowed R911012 to refurbish his holiday home.the loans require monthly repayments over 12 years. When he borrowed the money the interest rate was 12,4% per annum compounded monthly , but five years later the bank increased the annual interest rate to 13.9% in line with market rates. After 5 years the present value of the loan is R682 081,77 with new interest rate his monthly payments will increase by?

1
Expert's answer
2022-05-06T11:29:25-0400

If the present value of the loan is R911012 to be repaid after 12 years with an interest rate of 12.4% but to be changed to a new rate of 13.9% after the fifth year of starting, the monthly payments for the first five years is:"Pmt=\frac{PV}{[\frac{1-(1+\frac{r}{m})^{-(mn)}}{\frac{r}{m}}]}"

"Pmt=\frac{R911012}{[\frac{1-(1+\frac{12.4\%}{12})^{-(12×5)}}{\frac{12.4\%}{12}}]}"

"Pmt=R20450"


After the fifth year, the interest rate changes to 13.9%, the present value becomes R682018.77 and n becomes 7

"\therefore Pmt=\frac{R682081.77}{[\frac{1-(1+\frac{13.9\%}{12})^{-(12×7)}}{\frac{13.9\%}{12}}]}"

"Pmt=R12745"

His monthly payments decreases by R7705 (R20450-R12745).